write the equation y = -1/7 x - 3 in standard form. Thank you! :)

Algebra ->  Coordinate Systems and Linear Equations  -> Lessons  -> Linear Equations Lesson -> SOLUTION: Question 3: write the equation y = -1/7 x - 3 in standard form. Thank you! :)      Log On


   



Question 609059: Question 3:
write the equation y = -1/7 x - 3 in standard form.
Thank you! :)

Answer by ewatrrr(24785)

 
Hi
write the equation y = -1/7 x - 3 in standard form: Ax + By = C, where A > 0
y = -1/7 x - 3
y + (1/7)x = -3
(1/7)x + y = -3
x + 7y = -21